The base of a triangle is 2 m less than the height. The area of the triangle is 36m^2 . Determine the height of the triangle.

h=____

Please help I have no clue how to do this ,

Answer by scott8148(6628)

area equals 1/2 base times height

36=(1/2)(h-2)(h) __ multiplying by 2 __ 72=h^2-2h __ subtracting 72 __ 0=h^2-2h-72

use quadratic formula to find h __ h=[2 +- sqrt(4+288)]/2